# Feedback-Aware Monte Carlo Tree Search for Efficient Information Seeking in Goal-Oriented Conversations

### v1

@@ 5 Results and Discussion
[…]
The frequency of QGC for exhaustive tree search and MCTS differs significantly. An LLM call creates $m$ questions, each splitting $\Omega$ into affirmative and negative response sets, thereby giving rise to $2m$ nodes. For UoT, the first simulation requires $1+2m+(2m)^{2}+\dots+(2m)^{d_{s}}=\frac{(2m)^{d_{s}+1}-1}{2m-1}$ calls, and $(2m)^{d_{s}}$ calls for subsequent turns if paths are expanded till $d_{s}-1$, where $d_{s}$ is the depth-limit for simulation. For MCTS, the first expansion requires 1 call, and simulation for selecting a question at each turn requires $K\cdot d_{avg}$, where $K$ is the number of iterations and $0\leq d_{avg}\leq d_{s}$ because traversing requires 0 LLM calls when the nodes are already expanded. Exhaustive search scales exponentially with $d_{s}$, while MCTS scales linearly, highlighting its efficiency for deeper searches in the simulation phase.
Limitations. The system currently does not incorporate mechanisms to learn from mistakes in failure cases, which could potentially refine its decision-making process. Explicitly penalizing suboptimal or redundant questions also requires careful design of reward function and remains an open challenge. The framework can benefit from incorporating confidence metrics to better quantify uncertainty in risk-sensitive decision making domains. Addressing these limitations is a promising direction for future work.
6 Conclusion
In this paper, we introduced Monte Carlo Tree Search for Information Seeking Questions with Hierarchical Feedback (MISQ-HF), a novel framework for adaptive question-asking with a goal to optimize conversational efficiency and accuracy. By leveraging selective tree expansion and cluster-based feedback, our approach reduces computational overhead while maintaining high success rates across diverse domains such as medical diagnosis, troubleshooting, and general information-seeking tasks. We demonstrate the potential of combining structured and efficient planning with historical learning to dynamically adapt question-asking strategies. Our contributions aim to refine the decision-making process, enabling conversational AI systems to handle increasingly complex and open-ended scenarios with better planning.
Acknowledgments
This work is supported by the National Science Foundation (NSF) award IIS-2336769. The authors thank Hafsah Usman, student of the University of Washington, Seattle, for her valuable assistance in running a part of the experiments.
Appendix A Conversation Prompts for 20 Questions
A.1 Questioner Prologue
@@
A.2 Answerer/User Simulator Prologue
@@
If I guess X correctly in my question, directly respond “You guessed it. X is ‘{target_item}’.” instead of saying yes. Let us begin. Here is my first question.
@@ A.3 Targeting Prompt
A.4 Question Generation Prompt
@@
Appendix B Conversation Prompts for Medical Diagnosis
B.1 Questioner Prologue
You are a doctor, and your patient self-reports that: {problem_description}. {inform_set} You should ask your patient questions in English about symptoms which can only be answered by Yes or No, in order to find what disease this patient suffers from. Use the ongoing conversation for context to avoid redundant questions. Let us begin. Ask me the first question.
B.2 Answerer/User Prologue
You are the patient suffering from ‘{target_item}’, and I am the doctor. I will ask you up to 6 questions, and you should answer each one truthfully based on your disease, by saying Yes or No. Note that you must never reveal the disease until I tell it correctly. If I tell the disease correctly in my question, directly respond: “You are right. I am experiencing ‘{target_item}’.” Let us begin. Here is my first question.
B.3 Question Generation Prompt
@@ B.4 Targeting Prompt
The patient is probably suffering from one of the diseases below: {item_set_terminal}, or other from the initial set of possibilities.
Refer to the past conversation regarding the patient’s symptoms. Never repeat previously asked questions. Respond with 1 new question only, following the format: “Are you experiencing [disease name]?” Ensure that [disease name] was not asked before.
@@ B.5 Classify Prompt
{problem_description} For each disease under this report, if the patient is possibly affected, put this disease into ’YES: …’, otherwise to ’NO: …’. Your answer should strictly follow the template:
@@ B.6 Open Set Prompts
You are a doctor, and your patient self-reports that: {problem_description}. Please propose {size} diseases that you think your patient may suffer from. Your response should only be a list like: [”disease1”, ”disease2”, …]
@@
Based on the conversation history, please propose {size} diseases that your patient may suffer from. The list of {size} diseases should contain {existing_items}. Your response should only be a list like: [”disease1”, ”disease2”, …]
Appendix C Conversation Prompts for Troubleshooting
C.1 Questioner Prologue
@@
C.2 Answerer/User Prologue
You are the client with a device that has ‘{target_item}’, and I am the technician. I will ask you up to 20 questions, and you should answer each one truthfully based on the issue of your device, by saying Yes or No. Note that you must never reveal the issue name until I tell it correctly. If I tell your issue correctly in my question, directly respond: “You are right. My device has issues with ‘{target_item}’.” Let us begin. Here is my first question.
C.3 Question Generation Prompt
@@
Design a question to ask your client with a specific situation that can only be answered by YES or NO. Then classify the possible issues above based on this question. If the answer is ’YES’, put this issue into ’YES: …’, otherwise to ’NO: …’. Finally calculate how many issues are in YES and NO. Notably, this question should fulfill that the count of YES and NO are almost the same with a permissible discrepancy of no more than one!
@@
C.4 Targeting Prompt
Note that you should now point out and ask what issue the client is facing. Refer to the past conversation for context and never ask redundant questions. Respond with 1 new question only, following the format: “Are you experiencing [issue name]?” Ensure that [issue name] was not asked before.
@@ C.5 Classify Prompt
{problem_description} For each issue under this report, if the client is possibly facing it, put this issue into ’YES: …’, otherwise to ’NO: …’. Your answer should strictly follow the template:
@@ C.6 Open Set Prompts
You are a technician, and your client self-reports that: {problem_description}. Please propose {size} issues that you think your client may face with. Your response should only be a list like: [”issue1”, ”issue2”, …]
@@
Based on the conversation history, please propose {size} issues that your client may face with. The list of {size} issues should contain {existing_items}. Your response should only be a list like: [”issue1”, ”issue2”, …]
Appendix D Examples
Two sequences of history are maintained respectively for Questioner LLM and User (Answerer LLM). Each stores the questions asked and answers given so far. At the beginning of the conversation, the Question LLM is given the Questioner Prologue as the first prompt, and the Answerer LLM is given the Answerer Prologue as the first prompt. The rest of the conversation history remains the same, except for the switching between the “system” role and the “user” role for each LLM. Table 4, 5, 6 show example conversations from three domains.
S.No. System User
Problem Description:
Hello doctor, I have been in this situation for several years. Every summer, I often have diarrhea, but it has gotten worse in the past year or two. There is no regular pattern, sometimes it happens every two to three days when it’s mild, and two to three times a day when it’s severe. $\dots$ I went to a major hospital for examination before, and they found no parasite eggs or fungi, just a slight imbalance in the bacterial colonies, but the medication didn’t work. I don’t know what my condition is, I wonder if any doctor can provide some answers. Thank you. (Male, 28 years old).
1 Do you experience any symptoms related to the respiratory system, such as coughing or difficulty breathing? No, I don’t experience any symptoms related to the respiratory system, such as coughing or difficulty breathing. My main issue is with my stomach and bowel movements. I have been experiencing diarrhea, stomach aches, and urgent bowel movements, especially during the summer. The stool is often loose, sticky, and not well-formed, and it’s difficult to flush. I’ve been having this issue for several years, and it’s been getting worse over the past year or two.
2 Do you experience any abdominal pain or discomfort? Yes, I do experience abdominal pain and discomfort. It usually starts with my stomach aching more and more, and then I have to go to the bathroom. The pain can be quite severe, especially when the symptoms are bad. Sometimes, the stool is not very loose, but it’s accompanied by mucus and is not well-formed, which makes it difficult to flush. The abdominal pain and discomfort are quite frequent, and it’s been a persistent issue for me.
3 Do you experience any symptoms related to digestion, such as nausea or vomiting? I don’t usually experience nausea or vomiting, but I do have diarrhea, which can be quite loose and sticky at times. The stool is often not well-formed and can be difficult to flush. The frequency of the diarrhea can vary, sometimes it’s every two to three days when it’s mild, and sometimes it’s two to three times a day when it’s severe. The symptoms can be quite unpredictable and irregular, which makes it challenging to manage.
4 Are you experiencing symptoms of irritable bowel syndrome? You are right. I have Irritable bowel syndrome.

### v2

@@ Appendix B Open Set Results
Appendix C Conversation Prompts for 20 Questions
C.1 Questioner Prologue
@@
C.2 Answerer/User Simulator Prologue
@@
If I guess X correctly in my question, directly respond “You guessed it. X is ‘{target_item}’." instead of saying yes. Let us begin. Here is my first question.
@@ C.3 Targeting Prompt
C.4 Question Generation Prompt
@@
Appendix D Conversation Prompts for Medical Diagnosis
D.1 Questioner Prologue
You are a doctor, and your patient self-reports that: {problem_description}. {inform_set} You should ask your patient questions in English regarding their symptoms which can only be answered by Yes or No, in order to diagnose what disease this patient suffers. Carefully review the patient’s problem and the ongoing conversation. Avoid redundant questions. Let us begin. Ask me the first question.
The text “which can only be answered by Yes or No” is omitted in case of free-answer conversations.
D.2 Answerer/User Prologue
You are the patient suffering from ‘{target_item}’, and I am the doctor. I will ask you up to 6 questions, and you should answer each one truthfully based on your disease, by saying Yes or No. Note that you must never reveal the disease until I mention or ask about it. If I mention your disease in my question or ask about its symptoms, then you must directly respond “You are right. I am experiencing ‘{target_item}’." while saying Yes. Let us begin. Here is my first question.
In the beginning of the conversation:
You are a patient suffering from the disease of ‘{target_item}’, and communicating with a doctor. Here is your conversation history with another doctor: ‘{conv_history}
Remember the conversation above to answer current doctor’s question in English and do not reveal the disease until the doctor correctly mentions or asks about it. If the doctor mentions your disease in their question or asks whether you experience item, you must directly respond “You are right. I am experiencing ‘{target_item}’."
D.3 Question Generation Prompt
@@ D.4 Targeting Prompt
The patient is likely suffering from one of the diseases below: {item_set_terminal}, or other from the initial set of possibilities.
Refer the past conversation about patient’s symptoms. Never ask redundant questions. Respond with 1 new question only, following the format: “Are you experiencing [disease name]?" Ensure that [disease name] was not asked before.
@@ D.5 Classify Prompt
{problem_description} According to the patient’s symptoms, classify each possible disease given in the above set into ‘YES: …’ if even slightly possible, otherwise to ‘NO: …’ if completely unlikely. Only classify the diseases in above given set, do not add anything else. Your answer should be like:
@@ D.6 Open Set Prompts
You are a doctor, and your patient self-reports that: {problem_description}. Please propose {size} diseases that you think your patient may suffer from. Your response should only be a list like: [“disease1", “disease2", …]
@@
Based on the conversation history, please propose {size} diseases that your patient may suffer from. The list of {size} diseases should contain {existing_items}. Your response should only be a list like: [“disease1", “disease2", …]
Appendix E Conversation Prompts for Troubleshooting
E.1 Questioner Prologue
@@
E.2 Answerer/User Prologue
You are the client with a device that has ‘{target_item}’ and I am the technician. I will ask you up to 20 questions, and you should answer each one truthfully based on the issue of your device, by saying Yes or No. Note that you must never reveal the issue name until I tell it correctly. If I tell your issue correctly in my question, directly respond: “You are right. My device has issues with ‘{target_item}’." Let us begin. Here is my first question.
E.3 Question Generation Prompt
@@
Design a question to ask your client with a specific situation that can only be answered by YES or NO. Then classify the possible issues above based on this question. If the answer is ’YES’, put this issue into ‘YES: …’, otherwise to ‘NO: …’. Finally calculate how many issues are in YES and NO. Notably, this question should fulfill that the count of YES and NO are almost the same with a permissible discrepancy of no more than one!
@@
E.4 Targeting Prompt
Note that you should now point out and ask what issue the client is facing. The client is likely to be facing one of the issues below: {item_set_terminal}, so you must consider these. Refer the past conversation for problem context. Respond with 1 new question only, follow the format: “Are you experiencing [issue name]?" Ensure that [issue name] was not asked before.
@@ E.5 Classify Prompt
{problem_description} According to the user’s problem, classify each possible issue given in the above set into ’YES: …’if even slightly possible, otherwise to ’NO: …’if absolutely unlikely. Only classify the issues in above set, do not add any other. Your answer should be like:
@@ E.6 Open Set Prompts
You are a technician, and your client self-reports that: {problem_description}. Please propose {size} issues that you think your client may face with. Your response should only be a list like: [“issue1", “issue2", …]
@@
Based on the conversation history, please propose {size} issues that your client may face with. The list of {size} issues should contain {existing_items}. Your response should only be a list like: [“issue1", “issue2", …]
Appendix F Examples
Two sequences of history are maintained respectively for Questioner LLM and User (Answerer LLM). Each stores the questions asked and answers given so far. At the beginning of the conversation, the Question LLM is given the Questioner Prologue as the first prompt, and the Answerer LLM is given the Answerer Prologue as the first prompt. The rest of the conversation history remains the same, except for the switching between the “system" role and the “user" role for each LLM. Table 7, 6, 8 show example conversations from three domains.
@@ Appendix F Examples
S.No. System User
Problem Description:
At the age of 28, recently there has been a slight pain on both sides of the stomach, the stomach has been bothering for a few days, the stool is slightly bloody, and the stomach is slightly painful, but it comes in waves.
1 Do you experience symptoms that worsen after eating fatty foods? Actually, I’ve been experiencing a burning sensation in my stomach, and it seems to be triggered by certain foods and activities, including drinking alcohol. The pain is usually on both sides of my stomach and comes in waves. I’ve also had some bloody stools recently. However, I’m not sure if fatty foods specifically make it worse, but I do feel like my stomach is more bothered after drinking or eating certain things.
2 Do you experience abdominal pain that improves after eating? The pain in my stomach is a bit unpredictable, but it doesn’t really seem to improve after eating. In fact, sometimes I feel like my stomach is burning, and it can be uncomfortable. The pain is usually on both sides of my stomach and comes in waves. It’s been bothering me for about 20 days now, and I’ve also had some other symptoms like slightly bloody stools.
3 Are you experiencing a gastric ulcer? No, I’m not sure about that. I’ve been having some stomach issues, like pain on both sides of my stomach, and it’s been bothering me for about 20 days. I’ve also had some other symptoms like slightly bloody stools and a burning sensation in my stomach. But I haven’t been diagnosed with a gastric ulcer. I did talk to another doctor who suggested I take some medication, omeprazole, to help with my symptoms.
4 Are you experiencing gastritis? You are right. I have Gastritis.
@@
Appendix G QGC Efficiency
The frequency of QGC for exhaustive tree search and MCTS differs significantly. An LLM call creates $m$ questions, each splitting $\Omega$ into affirmative and negative response sets, thereby giving rise to $2m$ nodes. For UoT, the first simulation requires $1+2m+(2m)^{2}+\dots+(2m)^{d_{s}}=\frac{(2m)^{d_{s}+1}-1}{2m-1}$ calls, and $(2m)^{d_{s}}$ calls for subsequent turns if paths are expanded till $d_{s}-1$, where $d_{s}$ is the depth-limit for simulation. For MCTS, the first expansion requires 1 call, and simulation for selecting a question at each turn requires $K\cdot d_{avg}$, where $K$ is the number of iterations and $0\leq d_{avg}\leq d_{s}$ because traversing requires zero LLM calls when the nodes are already expanded. Exhaustive search scales exponentially with $d_{s}$, while MCTS scales linearly, highlighting its efficiency for deeper searches in the simulation phase.
Appendix H Broader Impacts
The proposed framework enhances goal-oriented conversational AI systems in domains like medical diagnosis and troubleshooting. By optimizing question-asking strategies, it improves decision-making success rates with reduced computational overhead. This can lead to more efficient diagnostic systems and customer support interactions, potentially benefiting both users and service providers. While general concerns about privacy and trustworthiness in LLMs persist, they are not specifically pertinent to the proposed system in this work. However, we acknowledge that use of LLMs as a core component of a system inherently brings challenges related to privacy and model biases, which must be carefully managed to ensure responsible deployment and maintain user trust.